OpenStudy (anonymous):

Functions/Trig help please! The piston in a lawn mower motor moves up and down 30 times in 1 s. Its total travel from the top position to the bottom position is 10cm. Vertical position y, in centimetres, versus time, t, in seconds can be modelled using a sinusoidal function. a) Write an equation to model the motion of the piston. (I have 5 for amplitude, which do I use Cos/Sin?, How do I get the compression/stretch factor?) b) Is this the only possible equation? Justify your answer.

OpenStudy (anonymous):

y=5cos(1/30*t)

OpenStudy (anonymous):

You could have a bunch of equations, because you could start the piston at t=0 at many different places ie top bottom middle

OpenStudy (anonymous):

The answer given is y = 5Sin(60*pi*t) . How did they get that?

OpenStudy (anonymous):

In particular, how'd they get 60*pi from the equation?

OpenStudy (anonymous):

30*2*pi from equation of omega or w

OpenStudy (anonymous):

w=2*pi*f f=(30 times per second) w=60*pi

OpenStudy (jdoe0001):

5 for the amplitude is correct, it does 30 cycles/periods in 1sec, so it takes 1/30 of a second to do 1 period keep in mind that regular period for sine will be \(\bf 2\pi\) so for the new period will be \(\bf \cfrac{1}{30}=\cfrac{2\pi}{\textit{new period}}\implies \textit{new period}=60\cdot 2\pi\)

OpenStudy (jdoe0001):

bah, more typos

OpenStudy (jdoe0001):

\(\bf \cfrac{1}{30}=\cfrac{2\pi}{\textit{new period}}\implies \textit{new period}=30\cdot 2\pi\)

OpenStudy (anonymous):

I haven't learned the equation of omega before... Clears things up for me now, thanks so much!

OpenStudy (jdoe0001):

and yes, you could use sine or cosine equally, depending on where the piston starts off

OpenStudy (jdoe0001):

\(\bf y = A \quad sin( \color{red}{B}x + C ) + D\\ \cfrac{1}{30}=\cfrac{2\pi}{\color{red}{\textit{B}}}\implies \color{red}{\textit{B}}=30\cdot 2\pi\)

OpenStudy (jdoe0001):

since we know the period is 1/30, I meant, thus the factor B will be that